# and cross estimates

### v1

@@ Abstract.
This work focuses on the propagation of chaos and the relaxation to Gibbs equilibrium for a system of $N$ classical Brownian particles with weak mean-field interactions. While it is known that propagation of chaos holds at rate $O(N^{-1})$ uniformly in time, and Gibbs relaxation at rate $O(e^{-ct})$ uniformly in $N$, we go a step further by showing that the cross error between chaos propagation and Gibbs relaxation is $O(N^{-1}e^{-ct})$. For translation-invariant systems on the torus, this leads to an improved mean-field approximation error at the level of the one-particle density: the error decreases from $O(N^{-1})$ to $O(N^{-1}e^{-ct})$. Our approach relies on a detailed analysis of the BBGKY hierarchy for correlation functions, and applies to both underdamped and overdamped Langevin dynamics with merely bounded interaction forces. We also derive new results on Gibbs relaxation and present partial extensions beyond the weak interaction regime.
@@ 1. Introduction
$\left\{\begin{array}[]{ll}\mathrm{d}X^{i,N}_{t}=V^{i,N}_{t}\,\mathrm{d}t,&t\geq 0,\\[2.84526pt] \mathrm{d}V^{i,N}_{t}=-\beta V^{i,N}_{t}\,\mathrm{d}t+\tfrac{\kappa}{N}\sum_{j=1}^{N}K(X^{i,N}_{t},X_{t}^{j,N})\,\mathrm{d}t-\nabla A(X_{t}^{i,N})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i},&t\geq 0,\\[5.69054pt] (X^{i,N}_{t},V^{i,N}_{t})|_{t=0}=(X^{i,N}_{\circ},V^{i,N}_{\circ})&\end{array}\right.$ (1.1)
@@
where $\beta>0$ is a given constant, and where $\kappa>0$ is a coupling parameter describing the strength of interactions. Throughout, we shall assume $\kappa$ to be small enough (independently of $N$) so as to ensure good enough properties of the mean-field dynamics — in particular, the uniqueness of the invariant measure. For the overdamped dynamics, some partial results can be obtained without this smallness condition; see Section 1.3 below.
@@
Note that this is not really restrictive in view of de Finetti’s theorem. We now consider a global weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}\cap\operatorname{L}^{1}(\mathbb{X}))$ of the Liouville equation (1.4) with these tensorized data. In the regime of a large number $N\gg 1$ of particles, a classical question then concerns the evolution of finite subsets of ‘typical’ particles in the system, as described by marginals of $F^{N}$,
@@
$F^{N,m}_{t}\,\to\,\mu_{t}^{\otimes m},\qquad\text{as\leavevmode\nobreak\ $N\uparrow\infty$},$ (1.7)
@@
where $\mathcal{W}_{2}$ is the 222-Wasserstein distance. Such a uniform-in-time estimate does not only rely on the propagation of initial chaos, but also requires some mechanism to ensure that correlations created by particle interactions are damped over time: this is typically a consequence of ergodic properties of the linearized mean-field operator, which indeed drives the correlation dynamics. On this matter, we also refer to [4, 3], where the damping of initial correlations is captured explicitly.
While propagation of chaos (1.9) amounts to estimating correlations in the particle system, it is natural to further investigate the size of higher-order correlations: in turn, this provides a hierarchy of higher-order corrections to the mean-field approximation (1.7) (see Bogolyubov correction in [15, 2] and higher orders in [23]). The two-particle correlation function is simply defined as
@@
up to some dependence on $m$. In recent years, such estimates have been established in various settings [15, 23], also uniformly in time [2, 33, 3].
An advantage of uniform-in-time estimates such as (1.9), or (1.11), is that they allow to connect mean-field approximation and relaxation to equilibrium; see Figure 1 below. Let us briefly recall what is known regarding relaxation to equilibrium, both at the level of the mean-field dynamics and of the $N$-particle system.
  1. For the mean-field dynamics (1.8), if $K$ is bounded and if $\kappa$ is small enough, it is well known that there is a unique steady state $M$ and that the solution $\mu_{t}$ converges exponentially,
@@
    where $Z[M]\in\mathbb{R}_{+}$ is the normalization constant ensuring $\int_{\mathbb{X}}M=1$. Note that this equation has indeed a unique solution provided $\kappa\beta\|W\|_{\operatorname{L}^{\infty}(\mathbb{R}^{d})}<1$.
  2. At the level of the $N$-particle system, for fixed $N$, the existence and uniqueness of the steady state $M^{N}$ for the Liouville equation (1.4) follows e.g. from [31] (without any smallness condition on $\kappa$). Although not explicit in general, it reduces to the usual Gibbs equilibrium in the case $K(x,y)=-\nabla W(x-y)$,
@@
[…]
@@
[…]
@@
    These estimates guarantee relaxation in relative entropy only for $t\gg\log N$, and in weighted $L^{2}$ norms only for $t\gg N$, in contrast with the $\mathcal{W}_{2}$-estimate (1.14) for $t\gg 1$. We will show in the sequel how to improve upon these results and remove the $N$-dependent prefactors. Also note that, to the best of our knowledge, no result was available on the uniform exponential relaxation when the interaction kernel $K$ is not Lipschitz — a case that we will also address here for the first time. Yet, the aforementioned works [4, 28, 20] have significantly relaxed the smallness condition on $\kappa$, which is not accessible to our methods (see however Section 1.3).
@@ 1.1. Main results
In this paper, we revisit the above questions regarding propagation of chaos, correlation estimates, and Gibbs relaxation. First, we show that, for $\kappa$ small enough, we can slightly improve previous results on those matters by a direct use of the BBGKY hierarchy, combined with bootstrap arguments as in [33] and with standard hypocoercivity estimates. We use $L^{2}$ norms with equilibrium weight $\omega\simeq M^{-1}$,
@@
[…]
@@
Beyond propagation of chaos and Gibbs relaxation, a natural question is whether we can have a relative error estimates for the relaxation of the mean-field error — or equivalently for the mean-field approximation of the $N$-particle relaxation error. In other words: can we show that the cross mean-field/relaxation error satisfies
$F_{t}^{N,m}-\mu_{t}^{\otimes m}-M^{N,m}+M^{\otimes m}\,=\,O(C^{m}N^{-1})\times O(C^{m}e^{-c_{0}t})\leavevmode\nobreak\ \text{?}$ (1.20)
This is our main result in this work, which we deduce from a hierarchical analysis taking advantage of uniform-in-time correlation estimates (1.17). A corresponding estimate is also obtained for the relaxation of correlation functions, cf. (1.22) below.
@@
[…]
@@
Corollary 1.3 (Improved error for translation-invariant mean field).
@@
1.2. Overdamped setting
@@
[…]
We turn to our main result on the cross mean-field/relaxation error: it can also be adapted directly and takes on the following guise.
Theorem 1.4.
Let $K\in L^{\infty}(\mathbb{R}^{2d})^{d}$, let $A$ satisfy (1.3), and let $\mu_{\circ}\in\mathcal{P}(\mathbb{R}^{d})\cap L^{2}(\omega)$. Consider the unique global weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}(\mathbb{R}^{dN})\cap\operatorname{L}^{2}(\omega^{\otimes N}))$ of the Liouville equation (1.24) with tensorized data (1.25). There exist $c_{0}>0$ (only depending on $d,K,A$) and $\kappa_{0},C>0$ (further depending on $\|\mu_{\circ}\|_{L^{2}(\omega)}$) such that given $\kappa\in[0,\kappa_{0}]$ we have for all $1\leq m\leq N$ and $t\geq 0$,
@@
Corollary 1.5 (Improved error for translation-invariant mean field).
Let $\mathbb{X}=\mathbb{T}^{d}$, $A\equiv 0$, and $K(x,y)=K_{0}(x-y)$ for some $K_{0}\in L^{\infty}(\mathbb{T}^{d})^{d}$. Given $\mu_{\circ}\in\mathcal{P}\cap L^{2}(\mathbb{T}^{d})$, consider the unique global weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}\cap L^{2}(\mathbb{T}^{dN}))$ of the corresponding Liouville equation (1.24) on $\mathbb{T}^{dN}$ with tensorized data (1.25), and the unique weak solution $\mu\in C(\mathbb{R}_{+};\mathcal{P}\cap\operatorname{L}^{2}(\mathbb{T}^{d}))$ of the corresponding mean-field equation (1.27) on $\mathbb{T}^{d}$. Then there exist $c_{0}>0$ (only depending on $d,K$) and $\kappa_{0},C>0$ (further depending on $\|\mu_{\circ}\|_{L^{2}(\mathbb{T}^{d})}$) such that given $\kappa\in[0,\kappa_{0}]$ we have for all $t\geq 0$,
@@
1.3. Beyond weak interactions
@@
[…]
@@
In Case 1, as $M^{N}=1$, the cross estimate (1.29) further implies for all $1\leq m\leq N$ and $t,\theta\geq 0$,

### v2

@@ Abstract.
We study the propagation of chaos and relaxation to Gibbs equilibrium for a system of $N$ classical Brownian particles with weak mean-field interactions. It is well known that propagation of chaos holds uniformly in time with rate $O(N^{-1})$ and that Gibbs relaxation holds uniformly in $N$ with exponential rate $O(e^{-ct})$. We go one step further by establishing a cross estimate that simultaneously captures both effects: the joint deviation between chaos propagation and Gibbs relaxation is of order $O(N^{-1}e^{-ct})$. In particular, for translation-invariant systems, this yields an accelerated propagation of chaos, with the mean-field approximation error at the level of the one-particle density improving from $O(N^{-1})$ to $O(N^{-1}e^{-ct})$. Our approach relies on a detailed analysis of the BBGKY hierarchy for correlation functions, and applies to both underdamped and overdamped Langevin dynamics with merely bounded interaction forces. In addition, we obtain new quantitative results on Gibbs relaxation and provide partial extensions beyond the weak interaction regime.
MSC-class: Primary 60K35; Secondary 35Q70, 82C22, 82C31, 35Q84, 91D30.
Keywords: interacting Brownian particles; propagation of chaos; Gibbs relaxation; cross estimates; BBGKY hierarchy; hypocoercivity.
@@ 1.1. General overview
[…]
@@
where $\beta>0$ is a given constant, and where $\kappa>0$ is a coupling parameter describing the strength of interactions. We shall primarily assume $\kappa$ to be small enough (independently of $N$) to guarantee good properties of the mean-field dynamics — in particular, the uniqueness of the invariant measure. For the overdamped dynamics, however, we also establish partial results that remain valid for arbitrary coupling strengths; see Section 1.4 below.
1.1.1. Propagation of chaos
@@
Note that this is not really restrictive in view of de Finetti’s theorem. We now consider a global weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}\cap{L}^{1}(\mathbb{X}))$ of the Liouville equation (1.4) with these tensorized data. In the regime of a large number $N\gg 1$ of particles, a classical question then concerns the evolution of finite subsets of ‘typical’ particles in the system, as described by marginals of $F^{N}$,
@@
$F^{N,m}_{t}\,\to\,\mu_{t}^{\otimes m},\qquad\text{as\penalty 10000\ $N\uparrow\infty$},$ (1.7)
@@
where $\mathcal{W}_{2}$ is the $2$-Wasserstein distance. Such a uniform-in-time estimate does not only rely on the propagation of initial chaos, but also requires some mechanism to ensure that correlations created by particle interactions are damped over time: this is typically a consequence of ergodic properties of the linearized mean-field operator, which indeed drives the correlation dynamics. On this matter, we also refer to [4, 3], where the damping of initial correlations is captured explicitly.
While propagation of chaos (1.9) amounts to estimating correlations in the particle system, it is natural to further investigate the size of higher-order correlations: in turn, this provides a hierarchy of higher-order corrections to the mean-field approximation (1.7) (see Bogolyubov correction in [15, 2, 16] and higher orders in [24]). The two-particle correlation function is simply defined as
@@
up to some dependence on $m$. In recent years, such estimates have been established in various settings [15, 24, 16], also uniformly in time [2, 34, 3].
1.1.2. Relaxation to equilibrium
An advantage of uniform-in-time estimates such as (1.9), or (1.11), is that they allow to connect mean-field approximation and relaxation to equilibrium; see Figure 1 below. Let us briefly recall what is known regarding relaxation.
At the level of the mean-field dynamics (1.8), if $K$ is bounded and if $\kappa$ is small enough, it is well known that there is a unique steady state $M$ and that the solution $\mu_{t}$ converges exponentially,
@@ 1.1.2. Relaxation to equilibrium
where $Z[M]\in\mathbb{R}_{+}$ is the normalization constant ensuring $\int_{\mathbb{X}}M=1$. Note that this equation has indeed a unique solution provided $\kappa\beta\|W\|_{{L}^{\infty}(\mathbb{R}^{d})}<1$.
At the level of the $N$-particle system, for fixed $N$, the existence and uniqueness of the steady state $M^{N}$ for the Liouville equation (1.4) follows e.g. from [32] (without any smallness condition on $\kappa$). Although not explicit in general, it reduces to the usual Gibbs equilibrium if $K(x,y)=-\nabla W(x-y)$,
@@
where $Z^{N}$ is the normalization constant. Similarly as for marginals of $F^{N}$, marginals of $M^{N}$ will be denoted by $\{M^{N,m}\}_{1\leq m\leq N}$. For $K$ Lipschitz and for $\kappa$ small enough, the uniform-in-$N$ exponential relaxation of the Liouville solution $F^{N}$ toward equilibrium $M^{N}$ was first established in [4, Remark 13] using coupling techniques: specifically, for all $1\leq m\leq N$ and $t\geq 0$,
@@
for some constants $C,c_{0}>0$ independent of $N,m$. Hypocoercivity techniques [33] ensure that the same must hold in $L^{2}$ norms with equilibrium weight, but with constants a priori depending on $N$. More recently, based on uniform-in-$N$ log-Sobolev and Poincaré inequalities, it was shown in [29, 21] that exponential relaxation in weighted $L^{2}$ norms and in relative entropy actually holds with a time constant $c_{0}$ independent of $N$, but a priori still with an $N$-dependent prefactor: more precisely, for all $1\leq m\leq N$ and $t\geq 0$,
@@
These estimates guarantee relaxation in relative entropy only for $t\gg\log N$, and in weighted $L^{2}$ norms only for $t\gg N$, in contrast with the $\mathcal{W}_{2}$-estimate (1.14) for $t\gg 1$. We will show in the sequel how to improve upon these results and remove the $N$-dependent prefactors.
1.1.3. Cross error estimate
The Bogolyubov correction to the mean-field description is the next-order contribution arising from pair correlations (see e.g. [15, 2, 16]), but its precise effects are often difficult to interpret. For conservative systems, the Lenard-Balescu theory predicts that this correction describes the very slow relaxation of the system toward equilibrium (see e.g. [15]). In the present setting of Brownian particles, however, diffusion already drives relaxation on an $O(1)$ timescale, which raises the question of what the corresponding Bogolyubov correction describes here. As we shall see, it simply accounts for the fact that the $N$-particle Gibbs equilibrium differs from the tensorized mean-field equilibrium: the correction captures the emergence of these nontrivial finite-$N$ equilibrium correlations in the long-time limit. This observation naturally leads us to studying the interplay between the mean-field approximation and the long-time relaxation: more precisely, our main result consists of the following ‘cross error’ estimate,
$F_{t}^{N,m}-\mu_{t}^{\otimes m}-M^{N,m}+M^{\otimes m}\,=\,O(C^{m}N^{-1})\times O(C^{m}e^{-c_{0}t}),$ (1.16)
which seems to be a new type of result in the field. This hybrid bound simultaneously recovers the optimal uniform-in-time mean-field estimate (1.9) and yields a uniform-in-$N$ version of the exponential relaxation bound (1.15). It thus quantifies how fast the finite-$N$ system approaches equilibrium while remaining close to the main-field description.
1.1.4. Overview of main results
In this work, we revisit several classical questions regarding propagation of chaos, correlation estimates, and relaxation to Gibbs equilibrium. Our analysis culminates in the proof of the cross error estimate (1.16); see Theorem 1.2 below. The proof combines a particularly detailed hierarchical analysis together with quantitative relaxation properties of the mean-field equation. Along the way, we also obtain a number of additional results of interest, both for Gibbs relaxation and for propagation of chaos:
  1. Uniform-in-NN Gibbs relaxation: For the Langevin dynamics (1.1), if the coupling constant $\kappa$ is small enough, we establish the first Gibbs relaxation estimate in $L^{2}$ that is fully uniform in $N$, thereby improving upon (1.15). This could be deduced from the cross error bound (1.16), but we rather derive it directly through a new hierarchical argument. To the best of our knowledge, this also provides the first exponential relaxation result beyond the case of a Lipschitz force kernel $K$.
  2. Time-accelerated propagation of chaos: For the corresponding translation-invariant Langevin dynamics on the torus, one has $M^{N,1}\equiv M$. The cross error estimate (1.16) then reduces to
    $F^{N,1}_{t}-\mu_{t}=O(N^{-1}e^{-c_{0}t}),$
    which yields an improved, time-accelerated rate for propagation of chaos on long timescales. To our knowledge, such a refined rate appears here for the first time in the literature.
  3. Strongly coupled setting: For the overdamped Langevin dynamics on the torus, we identify several situations in which the smallness assumption on the coupling constant $\kappa$ can be partially removed; see Section 1.4. This includes, in particular, the case when $K$ is translation-invariant and conservative. In that setting, a slightly more delicate variant of our argument yields the same cross error estimate up to an infinitesimal corrections in $N$,
    $F_{t}^{N,m}-\mu_{t}^{\otimes m}-M^{N,m}+M^{\otimes m}\,=\,O(C_{m}N^{-1}e^{-c_{0}t})+O(C_{m}N^{-\infty}).$
    As a consequence, we deduce a partial Gibbs relaxation result,
    $F^{N,m}-M^{N,m}\,=\,O(C_{m}e^{-c_{0}t})+O(C_{m}N^{-\infty}),$
    which ensures the expected exponential relaxation at least for $1\ll t\lesssim\log N$. Although not complete, this seems to be the only available relaxation result in this setting. In addition, we obtain the corresponding accelerated propagation of chaos, cf. Remark 1.7,
    $F^{N,m}-\mu_{t}^{\otimes m}\,=\,O(C_{m}N^{-1}e^{-c_{0}t})+O(C_{m}N^{-\infty}).$
Finally, we note that the present analysis can also be extended to Kac’s model. In a forthcoming work, we obtain analogous cross error estimates in that setting, together with a new uniform-in-$N$ exponential relaxation result.
@@ 1.2. Main results
Before turning to our main results on the cross error, we first revisit classical questions concerning propagation of chaos, correlation estimates, and relaxation to Gibbs equilibrium. For a sufficiently small coupling constant $\kappa$, we show that the earlier results on these matters can be slightly strengthened through a direct analysis of the BBGKY hierarchy, combined with bootstrap arguments as in [34] and standard hypocoercivity estimates. Throughout, we use weighted $L^{2}$ norms with equilibrium weight $\omega\simeq M^{-1}$,
@@
More precisely, our first proposition establishes two key estimates: (i) a uniform-in-time bound on correlation functions, and (ii) a uniform-in-$N$ Gibbs relaxation result. Item (i) extends the recent work [34] on the overdamped dynamics to the present underdamped setting, the adaptation relying only on hypocoercivity arguments. Item (ii) improves on the earlier work on the topic [4, 29, 21] for small $\kappa$: it requires merely bounded interaction forces $K$ and yields exponential relaxation in $L^{2}$ uniformly in $N$, thereby removing the $N$-dependent prefactor in (1.15). The proof, postponed to Section 2, relies on a new, self-contained hierarchical argument which, to the best of our knowledge, constitutes the first hierarchical approach to Gibbs relaxation. We note, however, that the aforementioned works [4, 29, 21] have managed to significantly relax the smallness condition on $\kappa$ when $K$ is Lipschitz, a feature that currently seems out of reach for our PDE-based method. In the overdamped case, a partial extension beyond the weak-interaction regime will be discussed in Section 1.4.
@@
We now turn to the interplay between the mean-field approximation and the long-time relaxation, and give a precise statement of the announced cross error estimate (1.16).
@@
$F^{N,m}_{t}$$\mu_{t}^{\otimes m}$$M^{N,m}$$M^{\otimes m}$rate $O(C^{m}N^{-1})$mean field $N\uparrow\inftyrate$ $O(e^{-ct})$relaxation $t\uparrow\inftyrate$ $O(e^{-ct})$rate $O(C^{m}N^{-1})$
@@
Corollary 1.3 (Accelerated propagation of chaos for translation-invariant systems).
@@
1.3. Extension 1: overdamped Langevin dynamics
@@
For the mean-field equation (1.27), uniqueness of the steady state $M$ follows from standard parabolic theory for $K$ bounded and $\kappa$ small enough, and the convergence is exponential; see e.g. Lemma 4.1 below. At the level of the $N$-particle system, the existence and uniqueness of the steady state $M^{N}$ for the Liouville equation (1.24) follows e.g. from [25, Section 3.3]. For $K$ Lipschitz and $\kappa$ small enough, as in the underdamped setting (1.14), the coupling argument in [4] ensures the uniform-in-$N$ exponential relaxation of the Liouville solution $F^{N}$ in $2$-Wasserstein distance. Convergence in relative entropy follows from the uniform-in-$N$ log-Sobolev inequality in [30]. More recently, in [22], the smallness condition on $\kappa$ was substantially weakened and exponential convergence in $L^{2}$ was deduced from a uniform-in-$N$ Poincaré inequality. However, as in the underdamped setting (1.15), estimates in relative entropy and in $L^{2}$ were obtained only up to $N$-dependent prefactors and they all required $K$ to be Lipschitz. For $\kappa$ small enough, the proof of Proposition 1.1(ii) is easily adapted to the present overdamped setting and thus again yields a partial improvement of previous work.
We turn to our main result on the cross error: it can also be adapted directly and takes on the following guise.
Theorem 1.4 (Cross mean-field/relaxation error).
Let $K\in L^{\infty}(\mathbb{R}^{2d})^{d}$, let $A$ satisfy (1.3), and let $\mu_{\circ}\in\mathcal{P}(\mathbb{R}^{d})\cap L^{2}(\omega)$. Consider the unique global weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}(\mathbb{R}^{dN})\cap{L}^{2}(\omega^{\otimes N}))$ of the Liouville equation (1.24) with tensorized data (1.25). There exist $c_{0}>0$ (only depending on $d,K,A$) and $\kappa_{0},C>0$ (further depending on $\|\mu_{\circ}\|_{L^{2}(\omega)}$) such that given $\kappa\in[0,\kappa_{0}]$ we have for all $1\leq m\leq N$ and $t\geq 0$,
@@
Corollary 1.5 (Accelerated propagation of chaos for translation-invariant systems).
Let $\mathbb{X}=\mathbb{T}^{d}$, $A\equiv 0$, and $K(x,y)=K_{0}(x-y)$ for some $K_{0}\in L^{\infty}(\mathbb{T}^{d})^{d}$. Given $\mu_{\circ}\in\mathcal{P}\cap L^{2}(\mathbb{T}^{d})$, consider the unique global weak solution $F^{N}\in C(\mathbb{R}_{+};\mathcal{P}\cap L^{2}(\mathbb{T}^{dN}))$ of the corresponding Liouville equation (1.24) on $\mathbb{T}^{dN}$ with tensorized data (1.25), and the unique weak solution $\mu\in C(\mathbb{R}_{+};\mathcal{P}\cap{L}^{2}(\mathbb{T}^{d}))$ of the corresponding mean-field equation (1.27) on $\mathbb{T}^{d}$. Then there exist $c_{0}>0$ (only depending on $d,K$) and $\kappa_{0},C>0$ (further depending on $\|\mu_{\circ}\|_{L^{2}(\mathbb{T}^{d})}$) such that given $\kappa\in[0,\kappa_{0}]$ we have for all $t\geq 0$,
@@
1.4. Extension 2: beyond the weak interaction regime
@@
In both cases, it can be shown that the mean-field steady state is unique and coincides with the Lebesgue measure for any $\kappa\geq 0$, and ergodic properties of linearized mean field were established in [12, Sections 3.6.2–3.6.3] (see in particular [8] for Case 2). A suitable adaptation of our analysis then leads to the following cross error estimates in weak norms up to $O(N^{-\infty})$. This implies in particular a uniform-in-$N$ Gibbs relaxation result up to $O(N^{-\infty})$, cf. (1.30) below, which is, to our knowledge, the first such result for those systems. The proof is postponed to Section 5.
@@
In Case 1, as $M^{N}=1$, the cross estimate (1.29) further implies the following accelerated propagation of chaos, for all $1\leq m\leq N$ and $t,\theta\geq 0$,